Find the extreme values of




x^4+y^4-2(x-y)^2

1
Expert's answer
2022-01-17T16:50:07-0500

"f(x,y)=x^4+y^4-2(x-y)^2\\
\dfrac{\partial f}{\partial x}=4x^3-4(x-y)=4(x^3-x+y)\\
\dfrac{\partial f}{\partial y}=4(y^3-y+x)\\
\text{At extremum, }\dfrac{\partial f}{\partial x}=\dfrac{\partial f}{\partial y}=0\\
\therefore x^3-x+y=0~~~(i), y^3-y+x=0~~~~(ii)\\
\text{Adding (i) and (ii), }\\
x^3+y^3=0\Rightarrow x=-y\\
\text{Putting x=-y in (i) and (ii), }\\
(x,y)=(0,0), (\sqrt 2, -\sqrt 2), (-\sqrt 2, \sqrt 2)\\
\dfrac{\partial^2 f}{\partial x^2}=4(3x^2-1), \dfrac{\partial^2 f}{\partial x^2}=4(3y^2-1)\\
\dfrac{\partial^2 f}{\partial x \partial y}=4=\dfrac{\partial^2 f}{\partial y \partial x}\\
\therefore \text{Hassein Matrix,} \\
H(x,y)=\begin{bmatrix} 12x^2-4&4\\4&12y^2-4\end{bmatrix}\\"


"\text{Determinant, } D(x,y)=144x^2y^2-48(x^2+y^2)\\
\therefore D(\sqrt2, -\sqrt2)=D(-\sqrt2, \sqrt2)=384>0\\
\dfrac{\partial^2 f}{\partial x^2}|_{(\sqrt2, -\sqrt2)}=\dfrac{\partial^2 f}{\partial x^2}|_{(-\sqrt2, \sqrt2)}
=20>0\\
\text{Hence, $f$ has two local minima at }\\
\text{$(\sqrt2, -\sqrt2)$ and $(-\sqrt2, \sqrt2)$}\\
~\\
\text{Now, } D(0,0)=0\\
\text{So, we can't decide anything by second derivative test}\\
\text{Now if we approach at (0,0) by $y=x$,}\\
f(x,y)=f(x,x)=2x^4>0\\
\text{Now if we approach at (0,0) by $y=-x$,}\\
f(x,y)=f(x,x)=2x^4-8x^2<0~~ \text{ near zero.}\\~\\
\text{So there is no extremum at $(0,0)$.}"